Butler County recorded 8 3/4 inches of rain during the past 5 months.
If this trend continues, how many inches of rain will fall in the next 2 months?

Answer by checkley75(3666)

8.75/5=x/2 cross multiply
5x=8.75*2
5x=17.5
x=17.5/5
x=3.5 inches of rain is expected in the next 2 months at the current trend.